OpenStudy (anonymous):

find the length of arc,S, on a radius 3m intercepted by a central angle of 210 degrees

OpenStudy (mertsj):

a central angle of 210 represents this part of the entire circle: 210/360=21/36=7/12 Find the circumference of the circle and multiply it by 7/12. That will be the length of the arc.

OpenStudy (anonymous):

could you please help me with another one? Find the area of the sector of the circle with radius 2 meters and central angle of (11pi)/6

OpenStudy (mertsj):

It's the same idea. A central angle of (11pi)/6 represents this part of the entire circle: [(11pi)/6]/(2pi)= 11/12 of the entire circle. Find the area of the circle and multiply it by 11/12. That will be the area of the sector.

OpenStudy (anonymous):

Thanks!

OpenStudy (mertsj):

yw